OpenStudy (anonymous):

Which of the following represents x^2 - 25/ x^2- 5x in simplified form

OpenStudy (btaylor):

factor the numerator and denominator, and cancel common factors:\[\frac{(x+5)(x-5)}{x(x-5)} = \frac{x+5}{x}\]